OpenStudy (anonymous):

Solve for x: 1 over 3 (2x − 8) = 4.

Nnesha (nnesha):

\[\huge\rm \frac{ 1 }{ 3 }(2x-8)=4\] like this ?

OpenStudy (anonymous):

Yes, sorry

Nnesha (nnesha):

alright what would be your first step ? any idea ?

OpenStudy (anonymous):

I have no idea. I think if you get me started, I might be able to

Nnesha (nnesha):

alright so to solve for x we should move all numbers to the right side remember the PEMDAS rule Parentheses Exponnets Multiply divide Add Subtract there is a fraction so start with the multiplication multiply both sides by 3

OpenStudy (anonymous):

1(2x - 8) = 16?

Nnesha (nnesha):

 let me know what you get after this